# Symplectic Quantization: numerical results for the Feynman propagator on a 1+1 lattice and the theoretical relation with Quantum Field Theory

### v2

@@ II Symplectic Quantization: from dynamics to ensemble averages
so that in Fourier space the “kinetic” contribution coming from each degree of freedom to the total action amounts to /2Planck-constant-over-$\hbar/2$. The relation in Eq. (29) can be also applied to the discretized momenta usually considered for a numerical simulation on the lattice:
$\langle\pi^{*}(k_{i})\pi(k_{i})\rangle=\frac{\hbar}{2}\quad\quad\forall\leavevmode\nobreak\ i.$ (30)
@@
$|\pi(k_{i};\tau=0)|^{2}=\hbar\quad\quad\forall\leavevmode\nobreak\ i,$ (31)
@@
III Simulation details
@@
$\Gamma$ :$:\bigg{\{}x:x_{\mu}=an_{\mu}\,,\quad n_{\mu}=-\frac{M_{\mu}}{2},...,\frac{M_{\mu}}{2}\quad\mu=0,1\bigg{\}}.$ (32)
@@
$\mathbb{H}[\phi,\pi]$ $=\frac{1}{2}\sum_{x\in\Gamma}\bigg{[}\pi(x)^{2}-(-1)^{s}\frac{1}{a^{2}}\phi(x)\Delta^{(0)}\phi(x)$
$-\frac{1}{a^{2}}\phi(x)\Delta^{(1)}\phi(x)+m^{2}\phi^{2}(x)+\frac{\lambda}{4}\phi^{4}(x)\bigg{]},$ (35)
@@
$F[\phi(x)]=-\frac{\delta\mathbb{H}[\phi,\pi]}{\delta\phi(x)}=$
$=\frac{(-1)^{s}}{a^{2}}\Delta^{(0)}\phi(x)+\frac{1}{a^{2}}\Delta^{(1)}\phi(x)-m^{2}\phi(x)-\lambda\phi^{3}(x),$ (37)
@@
An important point for the study of this paper is the definition of boundary conditions. We used two different kinds of boundary conditions for the simulations. For all results on Euclidean lattice and for the study of dynamics stability with or without non-linear interaction on Minkowski lattice, discussed respectively in Sec. IV and in Sec. V, we have used standard periodic boundary conditions on the lattice. Differently, in Sec. VI, aimed at studying the free propagation of physical signals across the lattice we used fringe boundary conditions [18], introduced with the purpose of mimicking the existence of an infinite lattice outside the simulation grid. Fringe boundary conditions are realized considering a larger lattice, which we denote as $\Gamma_{f}$, where the subscript “f” is for fringe, which is composed by the original lattice $\Gamma$ plus several additional layer of points which we denote as $\Gamma_{\text{ext}}$, in such a way that the fringe lattice is $\Gamma_{f}=\Gamma+\Gamma_{\text{ext}}$. For the fringe lattice one also considers periodic boundary conditions, but the generalized Hamiltonian for points belonging to $\Gamma$ and to $\Gamma_{\text{ext}}$ is different. Namely, the fringe lattice is characterized by the Hamiltonian:
@@
$\mathbb{H}_{\text{ext}}[\pi,\phi]$ $=\frac{1}{2}\sum_{x\in\Gamma}\bigg{[}\pi(x)^{2}+m^{2}\phi^{2}(x)+\frac{\lambda}{4}\phi^{4}(x)$
$+\epsilon\left(\frac{1}{a^{2}}\phi(x)\Delta^{(0)}\phi(x)-\frac{1}{a^{2}}\phi(x)\Delta^{(1)}\phi(x)\right)\bigg{]},$ (40)
where the coefficient $\epsilon$ is very small, $\epsilon\ll 1$. This choice of boundary conditions allows us to have a free propagation of signals across the boundary layer of $\Gamma$, our true simulation lattice, but the signal is then strongly damped when going across $\Gamma_{\text{ext}}$, the “external” boundary layer before making sort of interference at the periodic boundaries at the border of $\Gamma_{\text{ext}}$. This choice of boundary conditions allows us not only to deal with an overall system which is still Hamiltonian (apart from small corrections scaling as $1/L$), but also to have quite satisfactory results for the study of the Feynman propagator, as shown in Sec. VI.
@@
$|\phi(k;0)|^{2}=0\quad\quad\forall\leavevmode\nobreak\ k$
$|\pi(k;0)|^{2}=1\quad\quad\forall\leavevmode\nobreak\ k.$ (41)
IV Euclidean propagator
@@
$\langle$ $\phi(x_{1}),\ldots,\phi(x_{n})\rangle=$
[…]
@@
$\hat{\phi}(k,\tau)=\frac{a^{2}}{2\pi}\sum_{x\in\Gamma}e^{-i(k_{0}x_{0}+k_{1}x_{1})}\leavevmode\nobreak\ \phi(x,\tau),$ (44)
@@
$G(k)$ $=\langle\hat{\phi}^{*}(k)\hat{\phi}(k)\rangle=$
$\frac{1}{\Delta\tau}\sum_{i=0}^{M}\phi^{*}(k,\tau_{0}+\delta\tau_{i})\phi(k,\tau_{0}+\delta\tau_{i}),$ (45)
where $\delta\tau_{i}=i\cdot\delta\tau$. For the Euclidean lattice we have studied the lattice dynamics with the parameters and initial conditions given at the end of Sec. III, considering, in addition, value of mass $m=3.0$ and nonlinearity coefficient $\lambda=0.001$: the numerical value of the propagator in Fourier space perfectly reproduces the expected dumbbell shape, as shown in Fig. [1].
@@
V Minkowski lattice: linear and non-linear theory
@@
$\ddot{\phi}(k,\tau)+\omega_{k}^{2}\leavevmode\nobreak\ \phi(k,\tau)=0,$ (46)
@@
$\phi(k,\tau)$ $=\phi(k,0)\cos(\omega_{k}\tau)+\frac{\dot{\phi}(k,0)}{\omega_{k}}\sin(\omega_{k}\tau)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ \leavevmode\nobreak\ \omega_{k}^{2}>0$
$\phi(k,\tau)$ $=\phi(k,0)\cosh(z_{k}\tau)+\frac{\dot{\phi}(k,0)}{z_{k}}\sinh(z_{k}\tau)\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ \leavevmode\nobreak\ \omega_{k}^{2}<0,$ (48)
@@
$\omega_{k}^{2}>0\leavevmode\nobreak\ \leavevmode\nobreak$ $\Longrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \phi(k,\tau)=\frac{\sin(\omega_{k}\tau)}{\omega_{k}}$
$\omega_{k}^{2}<0\leavevmode\nobreak\ \leavevmode\nobreak$ $\Longrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \phi(k,\tau)=\frac{\sinh(z_{k}\tau)}{z_{k}}.$ (51)
@@
$\mathbb{H}[\phi,\pi]=\frac{1}{2}\int d^{d}k\leavevmode\nobreak\ \left(|\pi(k)|^{2}+\omega_{k}^{2}\leavevmode\nobreak\ |\phi(k)|^{2}\right),$ (52)
@@
$\omega_{k}^{2}>0\leavevmode\nobreak\ \leavevmode\nobreak\ \Longrightarrow$
$\mathbb{H}[\phi(\tau),\pi(\tau)]=\frac{1}{2}\int d^{d}k\leavevmode\nobreak\ \left[\cos^{2}(\omega_{k}\tau)+\sin^{2}(\omega_{k}\tau)\right]$
$\omega_{k}^{2}<0\leavevmode\nobreak\ \leavevmode\nobreak\ \Longrightarrow$
$\mathbb{H}[\phi(\tau),\pi(\tau)]=\frac{1}{2}\int d^{d}k\leavevmode\nobreak\ \left[\cosh^{2}(z_{k}\tau)-\sinh^{2}(z_{k}\tau)\right].$ (53)
@@
Figure 2: Behaviour of the normalized energy $(E(\tau)-E_{0})/E_{0}$ vs $\tau$ for a scalar free theory ($\lambda=0$) with $m=1.0$, $a=1.0$, $M_{\mu}=128$, with initial conditions $\pi(k;0)=1$ and $\phi(k;0)=0$ for all $k$’s. Notice the exponential growth with $\tau$ which sets in after a short transient.
Figure 3: Behaviour of the normalized energy $(E(\tau)-E_{0})/E_{0}$ vs $\tau$ for a scalar theory with a small self-interaction term, $\lambda=0.001$, and with $m=1.0$, $a=1.0$, $M_{\mu}=128$, with initial conditions $\pi(k;0)=1$ and $\phi(k;0)=0$ for all $k$’s. Oscillations are of order $\delta t$.
At the same time, attempting to study numerically the free-field dynamics on Minkowski lattice, the leap-frog algorithm, which proceedes alternating the update of kinetic and potential energy, cannot handle the situation where the overall energy is conserved but the two contributions diverge. Eventually, due to the accumulation of numerical errors, total energy starts to diverge exponentially as well with elapsing time, see Fig. 2 and the discussion below.
Since in the present section we are just interested in the stability of the theory, irrespectively of a realistic study of signals propagation across the lattice, we have considered for simplicity periodic boundary conditions. Let us remark that these conditions would not be appropriate for a more realistic study of two-point correlation functions with Minkowski metric, since in this case we would like to probe the causal structure of space-time: a signal escaping from the lattice at $+ct$ cannot appear back at $-ct$. For a similar reason even fixed boundary conditions would not be appropriate.
Using periodic boundary conditions we have checked numerically that the symplectic quantization dynamics of a free scalar field suffers from the pathology which can be conjectured already from the exact solution: after a certain time the whole energy starts to grow exponentially with $\tau$. In Fig. [2] we present the results of simulations of the free-field with Minkowski metric, all the parameters declared at the end of Sec. III and $m=1.0$, showing a clear evidence of the exponential divergence with $\tau$. What seemed a good solution to both the conceptual and numerical shortcomings of the free theory has been to consider that the physically relevant theory is only the interacting one: physical fields are always in interactions and the “free-field theory” is just an approximation, with some internal inconsistencies which are revealed by the symplectic quantization approach. Let us for instance consider a potential of the kind
@@
$\ddot{\phi}(x,\tau)=$
$-\partial_{0}^{2}\phi(x,\tau)+\sum_{i=1}^{d}\partial_{i}^{2}\phi(x,\tau)-m\phi(x,\tau)-\lambda\phi^{3}(x,\tau).$ (55)
Clearly, due to the non-linear term in Eq. (55), it is not possible anymore to diagonalize the equations in Fourier space, so that both the sin/cos and the sinh/cosh solutions cannot be taken into account as a reference. Yet to be proven mathematically, the stability of Eq. (55) is a quite delicate problem, since in general for many Fourier components the equations are linearly unstable. The intuition suggests that for each point of space-time $x$ the cubic force acts as a restoring term which prevents the amplitude $\phi(x,\tau)$ to grow without bounds. This intuition has been confirmed, up to the accuracy of our analysis, from our numerical results. By using periodic boundary conditions, the parameters and initial conditions declared at the end of Sec. III, setting the non-linearity coefficient $\lambda=0.001$ we find that the energy is no more divergent. The system relaxes to a stationary state with oscillations of order $|E(t)-E_{0}|/E_{0}=\mathcal{O}(\delta\tau)$, as is shown in Fig. 5.
Having assessed the stability of the symplectic quantization dynamics in the presence of non-linear interactions and periodic boundary conditions, it is now time to consider, keeping the non-linearity switched on, the more physical case of fringe boundary conditions [18]. This procedure will allow us to sample numerically the Feynman propagator for small non-linearity, as will be discussed in the next section.
@@
Figure 5: Real part of the two-point correlation function Fourier spectrum $G(k_{0},k_{1})=\langle\phi^{*}(k_{0},k_{1})\phi(k_{0},k_{1})\rangle$ (Feynman propagator) for a $\lambda\phi^{4}$ theory in $1+1$ space-time dimensions. Top): theoretical value of the free propagator with lattice spacing $a=1.0$, lattice side $M_{\mu}=128$, mass $m=3.0$; Down): numerical value from the interacting theory with the same parameters and nonlinearity $\lambda=0.001$. Initial conditions are set to $\phi(k;0)=0$ and $\pi(k;0)=1$ for all $k$’s. For this choice of parameters there are no unstable modes, i.e. for all $k$’s we have $\omega_{k}^{2}>0$.
VI Feynman propagator: numerical results
In the previous section we have shown how the presence of non-linear interactions solves the instability problem of the linear theory, still keeping periodic boundary conditions. But periodic boundary conditions are clearly unphysical, because one of the directions of our lattice corresponds to $ct$, so that periodicity of the boundaries is clearly meaningless. We need to devise a strategy to mimic the free propagation of any kind of signal across the boundaries as if outside there was an infinitely large lattice. This strategy is provided by the use of fringe boundary conditions, introduced in Sec. III.

### v3

@@ 2 Symplectic Quantization: from dynamics to ensemble averages
so that in Fourier space the “kinetic” contribution coming from each degree of freedom to the total action amounts to $\hbar/2$. The relation in Eq. (29) can be also applied to the discretized momenta usually considered for a numerical simulation on the lattice:
$\langle\pi^{*}(k_{i})\pi(k_{i})\rangle=\frac{\hbar}{2}\quad\quad\forall~i.$ (30)
@@
$|\pi(k_{i};\tau=0)|^{2}=\hbar\quad\quad\forall~i,$ (31)
@@
Before moving to the presentation of simulation details and the exposition of numerical results let us make a final remark on the assumption that dynamical averages along the intrinsic time Hamiltonian dynamics, like the one of Eq. (27), yields indeed a reliable sampling of the microcanonical probability density in Eq. (18). The general attitude in order to claim the equivalence between dynamical and ensemble sampling of observables expectation values is to assume a certain degree of ergodicity/chaoticity of the microscopic dynamics, which in general has to be considered case by case. To this respect, beside recalling that the case of a interacting scalar field theory with non-linear quartic potential studied here is generally considered the one of a system where the dynamics has good mixing properties, let us remark that in the present approach we do no really regard ergodicity as an issue. More precisely, we follow the perspective outlined in a recent review from one of us on the foundations of statistical mechanics BGVZ25 where it is explained how the use of statistical ensembles to draw predictions on the expectation values of generic observables requires ergodicity only in a weak sense, i.e., the mixing properties of the microscopic dynamics are not really the point. As for “generic” observables one has to think about all the observables which depend on large number of the system’s degrees of freedom. Just to make an example/analogy, we expect the behaviour of the two-point correlation function in real space for a weakly interacting field theory to “thermalize well” under the Hamiltonian dynamics in the same way that a good thermalization is achieved for individual particles in a classical harmonic chain, even if the Fourier modes of the harmonic chain does not thermalize at all. For further speculations on this point see, e.g., CVG22 .
3 Simulation details
@@
$\Gamma$ $:\bigg\{x:x_{\mu}=an_{\mu}\,,\quad n_{\mu}=-\frac{M_{\mu}}{2},...,\frac{M_{\mu}}{2}\quad\mu=0,1\bigg\}.$ (32)
@@
$\mathbb{H}[\phi,\pi]=\frac{1}{2}\sum_{x\in\Gamma}\bigg[\pi(x)^{2}-(-1)^{s}\frac{1}{a^{2}}\phi(x)\Delta^{(0)}\phi(x)-\frac{1}{a^{2}}\phi(x)\Delta^{(1)}\phi(x)+m^{2}\phi^{2}(x)+\frac{\lambda}{4}\phi^{4}(x)\bigg],$ (35)
@@
$F[\phi(x)]$ $=-\frac{\delta\mathbb{H}[\phi,\pi]}{\delta\phi(x)}=\frac{(-1)^{s}}{a^{2}}\Delta^{(0)}\phi(x)+\frac{1}{a^{2}}\Delta^{(1)}\phi(x)-m^{2}\phi(x)-\lambda\phi^{3}(x),$ (37)
@@
An important point for the study of this paper is the definition of boundary conditions. We used two different kinds of boundary conditions for the simulations. For all results on Euclidean lattice and for the study of dynamics stability with or without non-linear interaction on Minkowski lattice, discussed respectively in Sec. 4 and in Sec. 5, we have used standard periodic boundary conditions on the lattice. Differently, in Sec. 6, aimed at studying the free propagation of physical signals across the lattice we used fringe boundary conditions NNH99 , introduced with the purpose of mimicking the existence of an infinite lattice outside the simulation grid. Fringe boundary conditions are realized considering a larger lattice, which we denote as $\Gamma_{f}$, where the subscript “f” is for fringe, which is composed by the original lattice $\Gamma$ plus several additional layer of points which we denote as $\Gamma_{\text{ext}}$, in such a way that the fringe lattice is $\Gamma_{f}=\Gamma+\Gamma_{\text{ext}}$. For the fringe lattice one also considers periodic boundary conditions, but the generalized Hamiltonian for points belonging to $\Gamma$ and to $\Gamma_{\text{ext}}$ is different. Namely, the fringe lattice is characterized by the Hamiltonian:
@@
$\mathbb{H}_{\text{ext}}[\pi,\phi]=\frac{1}{2}\sum_{x\in\Gamma}\bigg[\pi(x)^{2}+m^{2}\phi^{2}(x)+\frac{\lambda}{4}\phi^{4}(x)+\epsilon\left(\frac{1}{a^{2}}\phi(x)\Delta^{(0)}\phi(x)-\frac{1}{a^{2}}\phi(x)\Delta^{(1)}\phi(x)\right)\bigg],$ (40)
where the coefficient $\epsilon$ is very small, $\epsilon\ll 1$. This choice of boundary conditions allows us to have a free propagation of signals across the boundary layer of $\Gamma$, our true simulation lattice, but the signal is then strongly damped when going across $\Gamma_{\text{ext}}$, the “external” boundary layer before making sort of interference at the periodic boundaries at the border of $\Gamma_{\text{ext}}$. This choice of boundary conditions allows us not only to deal with an overall system which is still Hamiltonian (apart from small corrections scaling as $1/L$), but also to have quite satisfactory results for the study of the Feynman propagator, as shown in Sec. 6.
@@
$|\phi(k;0)|^{2}=0\quad\quad\forall~k$
$|\pi(k;0)|^{2}=1\quad\quad\forall~k.$ (41)
4 Euclidean propagator
@@
$\langle$ $\phi(x_{1}),\ldots,\phi(x_{n})\rangle=\lim_{\Delta\tau\rightarrow\infty}\frac{1}{\Delta\tau}\int_{\tau_{0}}^{\tau_{0}+\Delta\tau}d\tau~\phi_{\mathbb{H}}(x_{1},\tau)\ldots\phi_{\mathbb{H}}(x_{n},\tau),$ (42)
@@
$\hat{\phi}(k,\tau)=\frac{a^{2}}{2\pi}\sum_{x\in\Gamma}e^{-i(k_{0}x_{0}+k_{1}x_{1})}~\phi(x,\tau),$ (44)
@@
$G(k)$ $=\langle\hat{\phi}^{*}(k)\hat{\phi}(k)\rangle=\frac{1}{\Delta\tau}\sum_{i=0}^{M}\phi^{*}(k,\tau_{0}+\delta\tau_{i})\phi(k,\tau_{0}+\delta\tau_{i}),$ (45)
where $\delta\tau_{i}=i\cdot\delta\tau$. For the Euclidean lattice we have studied the lattice dynamics with the parameters and initial conditions given at the end of Sec. 3, considering, in addition, value of mass $m=3.0$ and nonlinearity coefficient $\lambda=0.001$: the numerical value of the propagator in Fourier space perfectly reproduces the expected dumbbell shape, as shown in Fig. [1].
@@
5 Minkowski lattice: linear and non-linear theory
@@
$\ddot{\phi}(k,\tau)+\omega_{k}^{2}~\phi(k,\tau)=0,$ (46)
@@
$\phi(k,\tau)$ $=\phi(k,0)\cos(\omega_{k}\tau)+\frac{\dot{\phi}(k,0)}{\omega_{k}}\sin(\omega_{k}\tau)~~~~\forall~~\omega_{k}^{2}>0$
$\phi(k,\tau)$ $=\phi(k,0)\cosh(z_{k}\tau)+\frac{\dot{\phi}(k,0)}{z_{k}}\sinh(z_{k}\tau)~~\forall~~\omega_{k}^{2}<0,$ (48)
@@
$\omega_{k}^{2}>0$~~ $\Longrightarrow~~\phi(k,\tau)=\frac{\sin(\omega_{k}\tau)}{\omega_{k}}$
$\omega_{k}^{2}<0$~~ $\Longrightarrow~~\phi(k,\tau)=\frac{\sinh(z_{k}\tau)}{z_{k}}.$ (51)
@@
$\mathbb{H}[\phi,\pi]=\frac{1}{2}\int d^{d}k~\left(|\pi(k)|^{2}+\omega_{k}^{2}~|\phi(k)|^{2}\right),$ (52)
@@
$\omega_{k}^{2}>0~~\Longrightarrow\mathbb{H}[\phi(\tau),\pi(\tau)]=\frac{1}{2}\int d^{d}k~\left[\cos^{2}(\omega_{k}\tau)+\sin^{2}(\omega_{k}\tau)\right]$
$\omega_{k}^{2}<0~~\Longrightarrow\mathbb{H}[\phi(\tau),\pi(\tau)]=\frac{1}{2}\int d^{d}k~\left[\cosh^{2}(z_{k}\tau)-\sinh^{2}(z_{k}\tau)\right].$ (53)
@@
(a) Behaviour of the normalized energy $(E(\tau)-E_{0})/E_{0}$ vs $\tau$ for a scalar free theory ($\lambda=0$) with $m=1.0$, $a=1.0$, $M_{\mu}=128$, with initial conditions $\pi(k;0)=1$ and $\phi(k;0)=0$ for all $k$’s. Notice the exponential growth with $\tau$ which sets in after a short transient.
(b) Behaviour of the normalized energy $(E(\tau)-E_{0})/E_{0}$ vs $\tau$ for a scalar theory with a small self-interaction term, $\lambda=0.001$, and with $m=1.0$, $a=1.0$, $M_{\mu}=128$, with initial conditions $\pi(k;0)=1$ and $\phi(k;0)=0$ for all $k$’s. Oscillations are of order $\delta t$.
Figure 2: Comparison of normalized energy behaviour for scalar free theory and scalar theory with self-interaction.
At the same time, attempting to study numerically the free-field dynamics on Minkowski lattice, the leap-frog algorithm, which proceedes alternating the update of kinetic and potential energy, cannot handle the situation where the overall energy is conserved but the two contributions diverge. Eventually, due to the accumulation of numerical errors, total energy starts to diverge exponentially as well with elapsing time, see Fig. 2(a) and the discussion below.
Since in the present section we are just interested in the stability of the theory, irrespectively of a realistic study of signals propagation across the lattice, we have considered for simplicity periodic boundary conditions. Let us remark that these conditions would not be appropriate for a more realistic study of two-point correlation functions with Minkowski metric, since in this case we would like to probe the causal structure of space-time: a signal escaping from the lattice at +ct+ct cannot appear back at ct-ct. For a similar reason even fixed boundary conditions would not be appropriate.
Using periodic boundary conditions we have checked numerically that the symplectic quantization dynamics of a free scalar field suffers from the pathology which can be conjectured already from the exact solution: after a certain time the whole energy starts to grow exponentially with $\tau$. In Fig. [2(a)] we present the results of simulations of the free-field with Minkowski metric, all the parameters declared at the end of Sec. 3 and $m=1.0$, showing a clear evidence of the exponential divergence with $\tau$. What seemed a good solution to both the conceptual and numerical shortcomings of the free theory has been to consider that the physically relevant theory is only the interacting one: physical fields are always in interactions and the “free-field theory” is just an approximation, with some internal inconsistencies which are revealed by the symplectic quantization approach. Let us for instance consider a potential of the kind
@@
$\ddot{\phi}(x,\tau)=-\partial_{0}^{2}\phi(x,\tau)+\sum_{i=1}^{d}\partial_{i}^{2}\phi(x,\tau)-m\phi(x,\tau)-\lambda\phi^{3}(x,\tau).$ (55)
Clearly, due to the non-linear term in Eq. (55), it is not possible anymore to diagonalize the equations in Fourier space, so that both the sin/cos and the sinh/cosh solutions cannot be taken into account as a reference. Yet to be proven mathematically, the stability of Eq. (55) is a quite delicate problem, since in general for many Fourier components the equations are linearly unstable. The intuition suggests that for each point of space-time $x$ the cubic force acts as a restoring term which prevents the amplitude $\phi(x,\tau)$ to grow without bounds. This intuition has been confirmed, up to the accuracy of our analysis, from our numerical results. By using periodic boundary conditions, the parameters and initial conditions declared at the end of Sec. 3, setting the non-linearity coefficient $\lambda=0.001$ we find that the energy is no more divergent. The system relaxes to a stationary state with oscillations of order $|E(t)-E_{0}|/E_{0}=\mathcal{O}(\delta\tau)$, as is shown in Fig. 4. Let us remark, anticipating some of the upcoming results, that the instability of the deterministic dynamics of Eq. (13) in the case of a free theory is perfectly consistent with the shape of the canonical probability density $P[\phi]\sim\exp[S[\phi]/\hbar]$ for which we proved equivalence with the microcanonical weight $\rho_{\text{micro}}[\phi(x)]\sim\delta\left(\mathcal{A}-\mathbb{H}[\phi,\pi]\right)$, as will be shown in Sec. 7: $P[\phi]$ turns out to be ill-defined for the action of a free scalar field.
Having assessed the stability of the symplectic quantization dynamics in the presence of non-linear interactions and periodic boundary conditions, it is now time to consider, keeping the non-linearity switched on, the more physical case of fringe boundary conditions NNH99 . This procedure will allow us to sample numerically the Feynman propagator for small non-linearity, as will be discussed in the next section.
@@
Figure 4: Real part of the two-point correlation function Fourier spectrum $G(k_{0},k_{1})=\langle\phi^{*}(k_{0},k_{1})\phi(k_{0},k_{1})\rangle$ (Feynman propagator) for a $\lambda\phi^{4}$ theory in $1+1$ space-time dimensions. Top: theoretical value of the free propagator with lattice spacing $a=1.0$, lattice side $M_{\mu}=128$, mass $m=3.0$; Bottom: numerical value from the interacting theory with the same parameters and nonlinearity $\lambda=0.001$. Initial conditions are set to $\phi(k;0)=0$ and $\pi(k;0)=1$ for all $k$’s. For this choice of parameters there are no unstable modes, i.e. for all $k$’s we have $\omega_{k}^{2}>0$.
6 Feynman propagator: numerical results
In the previous section we have shown how the presence of non-linear interactions solves the instability problem of the linear theory, still keeping periodic boundary conditions. But periodic boundary conditions are clearly unphysical, because one of the directions of our lattice corresponds to $ct$, so that periodicity of the boundaries is clearly meaningless. We need to devise a strategy to mimic the free propagation of any kind of signal across the boundaries as if outside there was an infinitely large lattice. This strategy is provided by the use of fringe boundary conditions, introduced in Sec. 3.